# Saman Motameda,1 Laura Culpb Kevin Swerskyb Priyank Jainib,$\dagger$ and Robert Geirhosb,$\dagger$ aINSAIT

### v2

@@
Figure 4: How well do current video generative models understand physical principles? Left. The Physics-IQ score is an aggregated measure across four individual metrics, normalized such that pairs of real videos that differ only by physical randomness score 100%. All evaluated models show a large gap, with the best model scoring 24.1%, indicating that physical understanding is severely limited. Right. In addition, the mean rank of models across all four metrics is shown here; the Spearman correlation between aggregated results on the left and mean rank on the right is high (-.$\text{-}.87,\emph{p}<.005- .$), thus aggregating to a single Physics-IQ score largely preserves model rankings.
@@
Figure 5: Relationship between visual realism and physical understanding. Left. A multimodal large language model (Gemini 1.5 Pro) is asked to identify the generated video among the real and the generated video for each scenario (MLLM score) in a two-alternative forced choice paradigm. Chance rate is 50%; lower scores indicate that the model finds it harder to tell apart generated from real videos (= better realism). Sora-generated videos are hardest to distinguish from real videos for the model, whereas Lumiere (multiframe) is easiest. Right. Do models that produce ‘realistic-looking’ videos (as assessed by the MLLM score) also score better in terms of physical understanding (as assessed via the Physics-IQ score)? This scatterplot with linear fit and 95% confidence interval as a shaded blue area shows that this is not the case: Visual realism is uncorrelated with physical understanding (Pearson’s r = - 0.46, p=.247 not significant). Note that the y axis on this plot is inverted for easier interpretation (up & right are best).
@@ Results
The goal of our Physics-IQ benchmark is to understand, and quantify, whether generative video models learn physical principles. Therefore, we test all eight models in our study on every scenario and for each camera position (left, center, right) in the benchmark dataset. These samples are visualized in Figure 1. We first report the aggregated Physics-IQ results across all metrics related to physical understanding ($\mathsf{Spatial}\text{-}\mathsf{IoU},\leavevmode\nobreak\ \mathsf{Spatiotemporal}\text{-}\mathsf{IoU},\leavevmode\nobreak\ \mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU},\leavevmode\nobreak\ \mathsf{MSE}$) in Figure 4. The main takeaway from the left part of this figure is that all the models show a massive gap to the physical variance baseline, with the best model scoring only 24.1% out of the possible 100.0%. As we mentioned in the previous section, each scenario was recorded twice (take 1 and take 2) to estimate the natural variability in real-world physical phenomena. This estimate is termed the physical variance; the figure is normalized such that pairs of real videos that differ only by physical randomness score 100.0%. The gap between model performance and real videos demonstrates a severe lack of physical understanding in current powerful video generative models. Across the different models, VideoPoet (multiframe) (13) ranks best; interestingly, VideoPoet is a causal model. For the two models that have both an image2video (i2v) and a version conditioned on multiple frames (multiframe), the multiframe variants do better than the i2v variants. This is to be expected given that in order to predict the future as we require models to do on our challenging Physics-IQ benchmark, having access to temporal information (as multiframe variants have) should help.
Zooming in on these overall results, Figure 6 breaks down model performance into different categories that include solid mechanics, fluid dynamics, optics, thermodynamics, and magnetism. While there is no category that can be considered “solved”, performance varies across categories, with some showing promising indications and differences across models. Interestingly, all models perform much better on $\mathsf{Spatial}\text{-}\mathsf{IoU}$, a metric that has the weakest requirement in the sense that it is only sensitive to where an action occurred, not whether it occurred at the right time (as $\mathsf{Spatiotemporal}\text{-}\mathsf{IoU}$ would track) or whether it had just the right amount of action (as measured by $\mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU}$). Furthermore, even a relatively simple metric like $\mathsf{MSE}$ shows a large gap between physically realistic videos and model-generated predictions.
The performance of each model on each individual metric is reported in Table 1. As expected based on the aggregated results, VideoPoet (multiframe) performs best on a majority of the metrics (3 out of 4), while Stable Video Diffusion performs best on $\mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU}$ followed by VideoPoet. The fact that Stable Video Diffusion does well in terms of producing action, while struggling with correct object placement and dynamics (as evidenced by weaker $\mathsf{Spatial}\text{-}\mathsf{IoU}$ and $\mathsf{Spatiotemporal}\text{-}\mathsf{IoU}$ scores) is also qualitatively evident in the generated videos which show more than usual movement, with the movement frequently being physically implausible and exaggerated. Similarly, Sora is among the best performing models for $\mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU}$ but performs weaker on other metrics. Qualitatively, the generated videos from Sora are often visually and artistically superior, but they also frequently include transition cuts—despite instructed not to change the camera perspective—which is penalized by several other metrics. We expect that if a future version of this model more closely follows the prompt (static camera perspective, no camera movement), its Physics-IQ score would improve substantially. Qualitatively, success and failure cases are visualized in Figure 7.
@@
Physical Variance 0.645 0.512 0.626 0.002 100.0
VideoPoet (multiframe) 0.245 0.143 0.054 0.010 24.1
Runway Gen 3 (i2v) 0.220 0.109 0.044 0.015 18.4
Lumiere (multiframe) 0.170 0.146 0.034 0.013 18.2
VideoPoet (i2v) 0.175 0.106 0.057 0.012 18.0
Lumiere (i2v) 0.138 0.165 0.024 0.016 17.1
Stable Video Diffusion (i2v) 0.139 0.054 0.088 0.021 13.5
Pika 1.0 (i2v) 0.151 0.034 0.026 0.014 9.5
Sora (i2v) 0.142 0.041 0.055 0.036 8.7
@@ Discussion
Do video models learn physical principles?
We investigated the question whether the ability of video generative models to generate realistic-looking videos also implies that they have acquired an understanding of the physical principles that govern reality. Our benchmark shows that this is not the case: all evaluated models currently lack a deep understanding of physics. Even the highest-scoring model, VideoPoet (multiframe), only achieves a score of 24.1, falling significantly short of the best possible score of 100.0 obtained from the physical variance baseline, which quantifies the natural variability observed between real-world videos. That said, our results don’t imply that future models cannot learn a better physical understanding through next frame prediction. It remains an open question whether simply scaling the same paradigm further might solve this, or whether alternative (and possibly more interactive) training schemes might be required. Given the success of scaling deep learning, we are optimistic that future-frame prediction alone could lead to a much better understanding: while successful prediction does not guarantee successful understanding, acquiring a better understanding should certainly help with successful prediction. Learning physics by predicting pixels may sound challenging, but language models are known to learn syntax and grammar from text prediction alone (60).
It may be worth pointing out that even though the models in our study often failed to generate physically plausible continuations, most current models were already successful on some scenarios. For example, the highest-ranking model, VideoPoet (multiframe), displayed remarkable physical understanding in certain scenarios—such as accurately simulating paint smearing on glass. In contrast, many lower-ranking models exhibited fundamental errors, such as physically implausible interactions (e.g., objects passing through other objects). Separate studies (e.g. 40) based on synthetic datasets have shown that given a large enough dataset size, video models are able to learn specific physical laws. We consider it likely that as models are trained on larger and more diverse corpora of videos, their understanding of real-world physics will continue to improve. We hope that quantifying physical understanding, as we aim to do by open-sourcing the physics-IQ benchmark, might facilitate tracking progress in this area.
@@
Popular metrics to test quality and measure realism of generated videos include PSNR (43), FVD (45), LPIPS (46) and SSIM (44). However, designing metrics that quantify physical understanding in generated videos is a challenging endeavor. We proposed a suite of metrics to measure the spatial, temporal and perceptual coherence of video models. While individually none of these metrics is perfect, the combined insights from these metrics and the Physics-IQ score that integrates a normalized score across these metrics provide a holistic assessment of the strengths and weaknesses of video models. That said, none of these metrics directly quantify any physical phenomena and instead serve as proxies. For instance, the MLLM metric provided a way to quantify how much generated videos ‘deceive’ a multimodal model. However, the metric is limited by the capability of the underlying multimodal large language model (MLLM) itself. In our analysis, we found that while the MLLM was frequently able to identify generated videos (except for videos generated by Sora), its explanations for the decision were often wrong. As another example, we observed that Stable Video Diffusion often generated videos with large amounts of hallucinations and implausible object motions; nonetheless, it was the best performing model on $\mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU}$ showing that no metric should be assessed in isolation. This is also evidenced by the fact that e.g. Runway Gen 3 did very well in terms of the spatial location of actions ($\mathsf{Spatial}\text{-}\mathsf{IoU}$) while scoring poorly on temporal consistency ($\mathsf{Spatiotemporal}\text{-}\mathsf{IoU}$).
@@
Stabble Video Diffusion (i2v) 8 1024×576

### v3

# understand physical principles?

@@
Figure 4: How well do current video generative models understand physical principles? Left. The Physics-IQ score is an aggregated measure across four individual metrics, normalized such that pairs of real videos that differ only by physical randomness score 100%. All evaluated models show a large gap, with the best model scoring 29.5%, indicating that physical understanding is severely limited. Right. In addition, the mean rank of models across all four metrics is shown here; the Spearman correlation between aggregated results on the left and mean rank on the right is high (-.$\text{-}.92,\emph{p}<.005- .$), thus aggregating to a single Physics-IQ score largely preserves model rankings.
@@
Figure 5: Relationship between visual realism and physical understanding. Left. A multimodal large language model (Gemini 1.5 Pro) is asked to identify the generated video among the real and the generated video for each scenario (MLLM score) in a two-alternative forced choice paradigm. Chance rate is 50%; lower scores indicate that the model finds it harder to tell apart generated from real videos (= better realism). Sora-generated videos are hardest to distinguish from real videos for the model, whereas Lumiere (multiframe) is easiest. Right. Do models that produce ‘realistic-looking’ videos (as assessed by the MLLM score) also score better in terms of physical understanding (as assessed via the Physics-IQ score)? This scatterplot with linear fit and 95% confidence interval as a shaded blue area shows that this is not the case: Visual realism is uncorrelated with physical understanding (Pearson’s r = - 0.46, p=.249 not significant). Note that the y axis on this plot is inverted for easier interpretation (up & right are best).
@@ Results
The goal of our Physics-IQ benchmark is to understand, and quantify, whether generative video models learn physical principles. Therefore, we test all eight models in our study on every scenario and for each camera position (left, center, right) in the benchmark dataset. These samples are visualized in Figure 1. We first report the aggregated Physics-IQ results across all metrics related to physical understanding ($\mathsf{Spatial}\text{-}\mathsf{IoU},\leavevmode\nobreak\ \mathsf{Spatiotemporal}\text{-}\mathsf{IoU},\leavevmode\nobreak\ \mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU},\leavevmode\nobreak\ \mathsf{MSE}$) in Figure 4. The main takeaway from the left part of this figure is that all the models show a massive gap to the physical variance baseline, with the best model scoring only 29.5% out of the possible 100.0%. As we mentioned in the previous section, each scenario was recorded twice (take 1 and take 2) to estimate the natural variability in real-world physical phenomena. This estimate is termed the physical variance; the figure is normalized such that pairs of real videos that differ only by physical randomness score 100.0%. The gap between model performance and real videos demonstrates a severe lack of physical understanding in current powerful video generative models. Across the different models, VideoPoet (multiframe) (13) ranks best; interestingly, VideoPoet is a causal model. For the two models that have both an image2video (i2v) and a version conditioned on multiple frames (multiframe), the multiframe variants do better than the i2v variants. This is to be expected given that in order to predict the future as we require models to do on our challenging Physics-IQ benchmark, having access to temporal information (as multiframe variants have) should help. Zooming in on these overall results, Figure 6 breaks down model performance into different categories that include solid mechanics, fluid dynamics, optics, thermodynamics, and magnetism. While there is no category that can be considered “solved”, performance varies across categories, with some showing promising indications and differences across models. Interestingly, all models perform much better on $\mathsf{Spatial}\text{-}\mathsf{IoU}$, a metric that has the weakest requirement in the sense that it is only sensitive to where an action occurred, not whether it occurred at the right time (as $\mathsf{Spatiotemporal}\text{-}\mathsf{IoU}$ would track) or whether it had just the right amount of action (as measured by $\mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU}$). Furthermore, even a relatively simple metric like $\mathsf{MSE}$ shows a large gap between physically realistic videos and model-generated predictions. The performance of each model on each individual metric is reported in Table 1. As expected based on the aggregated results, VideoPoet (multiframe) performs best on a majority of the metrics (3 out of 4). Qualitatively, the generated videos from Sora are often visually and artistically superior, but they also frequently include transition cuts—despite instructed not to change the camera perspective—which is penalized by several other metrics. We expect that if a future version of this model more closely follows the prompt (static camera perspective, no camera movement), its Physics-IQ score would improve substantially. Qualitatively, success and failure cases are visualized in Figure 7.
@@
Physical Variance 0.678 0.535 0.577 0.002 100.0
VideoPoet (multiframe) 0.204 0.164 0.137 0.010 29.5
Lumiere (multiframe) 0.170 0.155 0.093 0.013 23.0
Runway Gen 3 (i2v) 0.201 0.115 0.116 0.015 22.8
VideoPoet (i2v) 0.141 0.126 0.087 0.012 20.3
Lumiere (i2v) 0.113 0.173 0.061 0.016 19.0
Stable Video Diffusion (i2v) 0.132 0.076 0.073 0.021 14.8
Pika 1.0 (i2v) 0.140 0.041 0.078 0.014 13.0
Sora (i2v) 0.138 0.047 0.063 0.030 10.0
@@ Discussion
Do video models understand physical principles?
We investigated the question whether the ability of video generative models to generate realistic-looking videos also implies that they have acquired an understanding of the physical principles that govern reality. Our benchmark shows that this is not the case: all evaluated models currently lack a deep understanding of physics. Even the highest-scoring model, VideoPoet (multiframe), only achieves a score of 29.5, falling significantly short of the best possible score of 100.0 obtained from the physical variance baseline, which quantifies the natural variability observed between real-world videos. That said, our results don’t imply that future models cannot learn a better physical understanding through next frame prediction. It remains an open question whether simply scaling the same paradigm further might solve this, or whether alternative (and possibly more interactive) training schemes might be required. Given the success of scaling deep learning, we are optimistic that future-frame prediction alone could lead to a much better understanding: while successful prediction does not guarantee successful understanding, acquiring a better understanding should certainly help with successful prediction. Learning physics by predicting pixels may sound challenging, but language models are known to learn syntax and grammar from text prediction alone (60).
It may be worth pointing out that even though the models in our study often failed to generate physically plausible continuations, most current models were already successful on some scenarios. For example, the highest-ranking model, VideoPoet (multiframe), displayed remarkable physical understanding in certain scenarios—such as accurately simulating paint smearing on glass. In contrast, many lower-ranking models exhibited fundamental errors, such as physically implausible interactions (e.g., objects passing through other objects). A studey based on synthetic datasets (40) has shown that given a large enough dataset size, video models are able to learn specific physical laws. We consider it likely that as models are trained on larger and more diverse corpora of videos, their understanding of real-world physics will continue to improve. We hope that quantifying physical understanding, as we aim to do by open-sourcing the physics-IQ benchmark, might facilitate tracking progress in this area.
@@
Popular metrics to test quality and measure realism of generated videos include PSNR (43), FVD (45), LPIPS (46) and SSIM (44). However, designing metrics that quantify physical understanding in generated videos is a challenging endeavor. We proposed a suite of metrics to measure the spatial, temporal and perceptual coherence of video models. While individually none of these metrics is perfect, the combined insights from these metrics and the Physics-IQ score that integrates a normalized score across these metrics provide a holistic assessment of the strengths and weaknesses of video models. That said, none of these metrics directly quantify any physical phenomena and instead serve as proxies. For instance, the MLLM metric provided a way to quantify how much generated videos ‘deceive’ a multimodal model. However, the metric is limited by the capability of the underlying multimodal large language model (MLLM) itself. In our analysis, we found that while the MLLM was frequently able to identify generated videos (except for videos generated by Sora), its explanations for the decision were often wrong. As another example, we observed that Stable Video Diffusion often generated videos with large amounts of hallucinations and implausible object motions; nonetheless, its $\mathsf{Spatial}\text{-}\mathsf{IoU}$ score is in the same ballpark as Lumiere, Sora, Pika and VideoPoet (i2v) showing that no metric should be assessed in isolation. This is also evidenced by the fact that e.g. Runway Gen 3 did very well in terms of the spatial location of actions ($\mathsf{Spatial}\text{-}\mathsf{IoU}$) while scoring poorly on temporal consistency ($\mathsf{Spatiotemporal}\text{-}\mathsf{IoU}$).
@@
Stable Video Diffusion (i2v) 8 1024×576